# Inferring Pluggable Types with Machine Learning

### v1

@@ VI-B RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
We rebuilt the GTN model with random samples of the dataset at 10% increments (10%, 20%, …) and ran the evaluation to see how much data is required for good results. The number of warnings by percentage of data trained is shown in fig. 6. What we see is a classic diagram of the model improving with training data until 60%. Performance then worsens after 70% with overfitting. In our feasibility study, performance is maximized from training around 50% of the dataset: around 16,000 classes. This estimate gives us a recommendation to calibrate the dataset size for similar tasks.
VII Limitations and Threats to Validity
VII-A Threats to the Validity of Our Experiments
@@
One threat to our NullAway experiment (section VI-A) is that the quality of the human annotations that we used as ground truth is somewhat questionable: they do not remove all NullAway warnings on 3 of the projects, which may distort our results. Another threat is the use of proxies for human effort: precision/recall of human-written annotations and reduction in warning count may both be poor proxies for effort. For example, even though NullGTN finds most of the annotations a human would write, those it fails to find might be the most difficult or complex to write. Another threat comes from the choice of benchmarks: we use programs that have been, at some point, annotated by a human to typecheck with NullAway, so that we have ground truth for our recall computation. However, in a deployment of NullGTN, we expect that the program would never have been human-annotated. Programs which have never been annotated might be trickier for inference, since human annotators might also change code to simplify the process of annotation (and those changes might stay in the code, even if the annotations themselves are later removed). The use of NullAway itself is also a threat to external validity (generalizability): our approach may not be as effective at inferring annotations that are useful for some other nullness analysis, such as Meta’s NullSafe.
@@
Our approach may not generalize beyond Java and beyond nullability. In theory, our approach should be applicable on other languages beyond Java, but evaluating that is beyond the scope of this work. We expect that the NaP-AST construction procedure in section III-A and a FastGTN architecture will be applicable to other pluggable typecheckers that prove the absence of other sorts of problems in Java, if a dataset of type qualifiers for those checkers were available. However, we cannot test whether this is effective in practice, since as far as we are aware sufficient training data does not exist for typecheckers that target problems other than nullability: there is at least an order of magnitude more code annotated for nullability than for any other typechecker in open-source, and in section VI-B we showed that NullGTN required about half of all the data we collected before it became effective.
VII-B Other Limitations of Our Approach
@@
All machine-learning-based techniques share one serious limitation: there are no guarantees that the annotations that it infers are correct. In our experiments, NullGTN demonstrated close alignment with human-written ground truth annotations—with a recall of 89%. Though there is also no guarantee that human-written annotations are correct, we believe that NullGTN’s ability to hew so closely to developers’ annotations suggests that this problem will be minor in a practical deployment: whether a human or NullGTN writes an incorrect annotation, a sound typechecker will reject it.
VIII Related Work
@@
The most closely-related work to ours is a recent approach for “error-driven” inference of nullability annotations for pluggable typecheckers [26]. The key idea of this appraoch is to treat the typechecker as a “black box” and use its error messages to drive inference: for example, if the checker warns that a return expression is nullable but that a return type is declared as nonnull, this approach could infer that an @Nullable annotation should logically be placed on the return type. In experiments using NullAway, this approach was shown to reduce the number of warnings emitted by the checker by 69.5% (vs. our 60%); they did not measure precision/recall of ground-truth annotations. However, our techniques are different: they relies on the checker’s output, while we rely on where humans have written nullability annotations historically (i.e., the training data for our model). In methodology, the two approaches are orthogonal; we hope to profitably combine them in future work, since the differing approaches mean that they probably infer different subsets of nullability annotations. Another closely related approach bootstraps global inference from the local inference that is built into most pluggable typechecking frameworks [27]. This approach, though not specific to nullness, also uses a fundamentally different approach and might be profitably combined with ours in future work: it works deductively, unlike our approach.
@@
Some prior works [1, 28] use constraint-based type inference for pluggable types. These approaches do not support nullness (they are specific to a units-of-measurement type system), and cannot be readily applied to other pluggable type systems: each constraint-based inference approach is unique. Our methodology, by contrast, could be applied to any type system, if training data were available: a model developer would need to repeat our ablation experiments to drop appropriate nodes from the NaP-AST, but otherwise could reuse our infrastructure. Prior approaches for explaining type inference errors [29, 30, 31, 32] and for gradual type migration [33] suffer the same problems as other constraint-based appraoches (i.e., do not apply easily to new type systems). There is also prior work that involves the user in type inference [34], which is complementary to ours. The Daikon invariant detector [35] can insert @Nullable annotations at program locations that can be null at run time [36]. Like any dynamic approach, it is inherently limited by the quality of program’s test suite.
@@
There is a lot of work on inference of static types for dynamically-typed languages such as Python or JavaScript using machine learning [37, 38, 6, 7, 8, 9, 10, 11]. These approaches are not directly applicable to our problem, since they do not reason about nullability, but about the base type system, and their models and encodings are specialized to that problem.
@@
There is a large literature on type inference and type reconstruction algorithms based on constraint solving [39, Chapter 22]. These techniques build and then solve a system of constraints; the specific constraints are type-system-dependent, which makes this approach unsuitable for pluggable types: determining the constraints to solve would require translating the typechecker’s logic into the logic of the constraint solver, which effectively would require re-implementing the typechecker. Recent work [40] has proposed an extension to the classic constraint-based type inference approach Algorithm W [41] that enables reasoning about nullability. However, extending this to the mainstream object-oriented languages (and their associated nullability checkers) is an open-problem.

### v2

@@ 7.3. RQ2
Figure 8 shows the result: a classic diagram of the model improving with training data until 60%; performance then worsens after 70% with overfitting. Performance is maximized from training around 50% of the dataset: around 16,000 classes. This estimate gives us a recommendation to calibrate the dataset size for similar tasks.
7.3.1. Availability of Data for Other Pluggable Type Systems
We collected annotations using Sourcegraph’s (Slack and Liu, 2013) src tool to investigate whether any other pluggable type systems (for Java) besides nullability have enough publicly-available data to plausibly train them the same way we trained NullGTN. We found that none did. The closest were @GuardedBy (from the Checker Framework’s Lock Checker (Ernst et al., 2016)) with 1,752 annotated classes; and integer @NonNegative and @Positive annotations (from several sources, including the Checker Framework’s Index Checker (Kellogg et al., 2018)) with 1,675 and 1,008 annotated classes, respectively. These occur an order of magnitude less often than @Nullable, showing there is insufficient data for us to repeat our experiments with another pluggable type system until humans annotate more classes: a “chicken-and-egg problem” for type qualifier inference.
7.3.2. Discussion: “Chicken-and-Egg Problem” for Type Qualifier Inference
The present work does not try to solve this “chicken-and-egg problem” of needing data to enable type qualifier inference (as proposed in this paper), while human adoption is needed to produce that same data. Our contribution is to show that simple feature engineering and choice of model is sufficient to do a reasonable job: that is, we have shown the feasibility of our approach, if enough data is available. However, our work does give the community a sense of how much “human adoption” is necessary before training an inference model like NullGTN is practical, which we believe is a useful step towards an otherwise intractable problem. Further, our work shows that once the threshold for ML-based qualifier inference is reached, it can be leveraged to help increase adoption more.
8. Limitations and Threats to Validity
8.1. Threats to the Validity of Our Experiments
@@
One threat to our NullAway experiment (section 7.1) is that the quality of the human annotations that we used as ground truth is somewhat questionable: they do not remove all NullAway warnings on 3 of the projects, which may distort our results. Another threat is the use of proxies for human effort: precision/recall of human-written annotations and reduction in warning count may both be poor proxies for effort. For example, even though NullGTN finds most of the annotations a human would write, those it fails to find might be the most difficult or complex to write. Another threat comes from the choice of benchmarks: we use programs that have been, at some point, annotated by a human to typecheck with NullAway, so that we have ground truth for our recall computation. However, in a deployment of NullGTN, we expect that the program would never have been human-annotated. Programs which have never been annotated might be trickier for inference, since human annotators might also change code to simplify the process of annotation (and those changes might stay in the code, even if the annotations themselves are later removed). The use of NullAway itself is also a threat to generalizability: our approach may not be as effective at inferring annotations that are useful for some other nullness analysis, such as Meta’s NullSafe.
@@
Our approach may not generalize beyond Java or nullability. In theory, our approach could be applied to other languages beyond Java, but evaluating that is out-of-scope. We expect that the NaP-AST construction procedure in section 3.2 and a FastGTN architecture will be applicable to other pluggable typecheckers that prove the absence of other problems in Java, if a dataset of type qualifiers for those checkers were available. However, we cannot test whether this is effective in practice, since as we showed in section 7.3.1 sufficient training data does not exist for typecheckers that target problems other than nullability: there is at least an order of magnitude more code annotated for nullability than for any other typechecker in open-source, and in section 7.3 we showed that NullGTN required about half of all the data we collected before it became effective.
Another threat to generalizability is that the data experiment in section 7.3 may not be accurate for other pluggable type systems. In particular, nullability type systems, compared to other extant pluggable type systems, are both relatively simple (only two type qualifiers, in a relatively simple type hierarchy) and relatively annotation-intensive (because null is pervasive in most Java programs). We believe that both of these features probably bias nullability inference to require less data, so our estimate may be overly conservative—type systems with large type hierarchies, dependent types, or other complex features may require much more data than our estimates suggest.
8.2. Generality
Our approach works for pluggable type systems with a finite set of types, assuming sufficient training data. The restriction to a finite set of types is an important limitation: we must have fixed the set of type qualifiers to place before training the model, because we’re using a classifier. Many pluggable type systems, such as units-of-measure (Xiang et al., 2020) or integer signedness (e.g., @NonNegative and @Positive in section 7.3.1), fit this model. However, systems with parameterized qualifiers, like @GuardedBy (section 7.3.1), require a different approach—possibly combining our method (to predict annotations) with a generative model (to generate parameters).
In practice, human-annotated code is an order of magnitude rarer for other checkers than the amount needed to train NullGTN (shown in section 7.3.1). This prevents us from directly evaluating the generality of our approach. We’re investigating approaches to synthesize training data for other pluggable type systems, but that is another, separate problem.
8.3. Other Limitations of Our Approach
@@
All machine-learning-based techniques share one serious limitation: there are no guarantees that the annotations that it infers are correct. In our experiments, NullGTN demonstrated close alignment with human-written ground truth annotations—with a recall of 69%. Though there is also no guarantee that human-written annotations are correct, we believe that NullGTN’s ability to hew so closely to developers’ annotations suggests that this problem will be minor in a practical deployment: whether a human or NullGTN writes an incorrect annotation, a sound typechecker will reject it.
9. Related Work
@@
The most closely-related work to ours is a recent approach for “error-driven” inference of nullability annotations for pluggable typecheckers (Karimipour et al., 2023). The key idea of this approach is to treat the typechecker as a “black box” and use its error messages to drive inference: for example, if the checker warns that a return expression is nullable but that a return type is declared as nonnull, this approach could infer that an @Nullable annotation should logically be placed on the return type. In experiments using NullAway, this approach was shown to reduce the number of warnings emitted by the checker by 69.5% (vs. our 60%); they did not measure precision/recall of ground-truth annotations. However, our techniques are different: they relies on the checker’s output, while we rely on where humans have written nullability annotations historically (i.e., the training data for our model). In methodology, the two approaches are orthogonal; we hope to profitably combine them in future work, since the differing approaches mean that they probably infer different subsets of nullability annotations. Another closely related approach bootstraps global inference from the local inference that is built into most pluggable typechecking frameworks (Kellogg et al., 2023). This approach, though not specific to nullness, also uses a fundamentally different approach and might be profitably combined with ours in future work: it works deductively, unlike our approach.
@@
Some prior works (Xiang et al., 2020; Li, 2017) use constraint-based type inference for pluggable types. These approaches do not support nullness (they are specific to a units-of-measurement type system), and cannot be readily applied to other type systems: each is unique. Our methodology, by contrast, could be applied to any type system, if training data were available: a model developer would need to repeat our ablation experiments to drop appropriate nodes from the NaP-AST, but otherwise could reuse our infrastructure. Prior approaches for explaining type inference errors (Lerner et al., 2007; Pavlinovic et al., 2014; Loncaric et al., 2016; Zhang et al., 2017) and for gradual type migration (Phipps-Costin et al., 2021) suffer the same problems as other constraint-based approaches (i.e., do not apply easily to new type systems). There is also prior work that involves the user in type inference (Vakilian et al., 2015), which is complementary to ours. The Daikon invariant detector (Ernst et al., 2007) can insert @Nullable annotations at program locations that can be null at run time (The Daikon developers, 2023). Like any dynamic approach, it is inherently limited by the quality of program’s test suite.
@@
There is a lot of work on inference of static types for dynamically-typed languages such as Python or JavaScript using machine learning (Xu et al., 2016; Pradel et al., 2020; Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021). These approaches are not directly applicable to our problem, since they do not reason about nullability, but about the base type system, and their models and encodings are specialized to that problem.
@@
There is a large literature on type inference and type reconstruction algorithms based on constraint solving (Pierce, 2002, Chapter 22). These techniques build and then solve a system of constraints; the specific constraints are type-system-dependent, which makes this approach unsuitable for pluggable types: determining the constraints to solve would require translating the typechecker’s logic into the logic of the constraint solver, which effectively would require re-implementing the typechecker. Recent work (Madsen and van de Pol, 2021) has proposed an extension to a classic constraint-based type inference approach (Damas and Milner, 1982) that enables reasoning about nullability. However, extending this to the mainstream object-oriented languages (and their associated nullability checkers) is an open-problem.
10. Data Availability
The data and code from our experiments are open-source and available (Kazi Siddiqui, Martin Kellogg, [n. d.]).